# DMRetriever: A Family of Models for Improved Text Retrieval in Disaster Management

### v1

@@ 1 Introduction
Disasters, both natural and man-made, inevitably disrupt communities, endanger lives, and cause severe economic losses (Fan et al., 2021; Liu et al., 2025). During such events, emergency responders, policymakers, and affected populations rely on information retrieval (IR) systems for decision-making, situational awareness, and coordination of relief efforts (Abbas and Miller, 2025). Timely and reliable information access is thus essential for effective disaster management (Priya et al., 2020).
@@
Information needs during disaster events are highly diverse, spanning fact-checking critical information, monitoring social media, and consulting technical or policy documents (See §3.1 for full lists of search intents) (Yin et al., 2025). However, a recent analysis of 30 popular general-domain IR models finds that none consistently achieve SOTA performance across all disaster management-specific search intents (Yin et al., 2025). This limits their real-world applicability: some queries yield valuable insights, while others return irrelevant results, delaying decision-making and risking missed evidence that could compromise response efforts.
@@
Figure 2: Overview of data refinement pipeline and three-stage training framework for DMRetriever. MTP: Massive Text Pairs, MTT: Massive Text Triplets. Both MTP and MTT in sub-figure (B) include DM (Disaster Management) and GD (General Domain) variants, with constructions of GD variants omitted here and detailed in §3.3.1 and §3.4.3. DM passages include those extracted from PDFs and those generated by LLMs. DMRetriever-Init: backbone after bidirectional attention adaptation for decoder-only; original backbone for encoder-only. DMRetriever-PT: model after pre-training. DMRetriever-It2: model after iteration 2. Algorithm 1 details the full workflow.
@@
The success of domain-specific IR models in fields such as biomedicine (Xu et al., 2024b), law (Chalkidis et al., 2020), and finance (Sarmah et al., 2024) shows the effectiveness of adding deep domain knowledge to improve performance. However, injecting disaster management knowledge into DMRetriever to improve performance poses two challenges: (1) the lack of high-quality, knowledge-rich data that covers diverse search intents during disasters, and (2) the absence of an efficient training method that can transfer such knowledge across models of different sizes.
To tackle the first challenge, we propose a data refinement pipeline (A in Figure 2) to generate the first, knowledge-rich data for disaster management. We first prompt a large language model (LLM) [^1] We use GPT-4o-mini as LLM in this work. with domain-specific passages to generate large-scale unlabeled query–passage pairs [^2] Following Xiao et al. (2024); Wang et al. (2022), we treat data without mined negatives as unlabeled and data with mined hard negatives as labeled. (A-1 in Figure 2, §3.3.1). Given limitations of existing IR models in disaster management, we refine these noisy corpora with mutual-agreement filtering, which removes mislabeled pairs through model consensus (A-2 in Figure 2, §3.4.2). Recognizing the varied learning capacities of different-sized DMRetriever models, we further propose difficulty-aware hard negative mining to generate negatives at varying difficulty levels (A-3 in Figure 2, §3.4.3), finally yielding a high-quality labeled dataset.
For the second challenge, we propose a three-stage training framework (B in Figure 2) that enables DMRetriever models of different sizes to effectively learn domain knowledge. For large variants initialized from decoder-only backbones (Table 1), we adapt their causal attention into bidirectional attention, allowing access to information from future tokens (B-1 in Figure 2, §3.2). Next, to better initialize DMRetriever and inject domain knowledge, we perform large-scale contrastive pre-training on unlabeled query–passage pairs (B-2 in Figure 2, §3.3). Finally, to adapt DMRetriever to diverse search intents, we perform instruction fine-tuning on labeled data2, where for small variants that have limited learning capacity, the difficulty of negative samples is progressively increased during fine-tuning (B-3 in Figure 2, §3.4.4).
@@
    We design a three-stage training framework enabling models of varied sizes to effectively absorb domain knowledge and enhancing their ability across diverse search intents.
@@ 2 Related Work
General-domain retrieval has been long dominated by bidirectional encoders such as BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020), with models like Sentence-BERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) improving performance through contrastive pre-training and supervised fine-tuning on MS MARCO (Nguyen et al., 2016). More recently, decoder-only models, from GPT-based embeddings (Brown et al., 2020) to models such as Qwen3-Embedding (Zhang et al., 2025), have leveraged instruction tuning and large-scale synthetic data to achieve SOTA results.
However, these general-domain models face clear limitations in disaster management, with no approach achieving SOTA performance across diverse search intents (Yin et al., 2025) due to distribution shift issue (Thakur et al., 2021). Prior work on disaster information access has mainly targeted on social media, emphasizing clustering of posts, event or situational classification, and named entity recognition (Imran et al., 2015; Alam et al., 2021; Yin et al., 2024). While valuable, these efforts remain narrow in scope and have yet to address the need for robust retrieval systems tailored to disaster management-related diver search intents.
@@ 3.1 Preliminaries
In DMRetriever, queries and passages are encoded into dense vectors using a shared embedding model $\mathbf{E}(\cdot)$. Given a query $q$ and a passage $p$, we prepend a search intent-specific instruction $I_{q}$ to the query to improve search intent awareness (Wei et al., 2021; Ouyang et al., 2022; Lee et al., 2024a).
Following DisastIR (Yin et al., 2025), we consider six disaster management-specific search intents: question-answer (QA), Twitter (TW), Fact Checking (FC), Natural Language Inference (NLI), and Semantic Textual Similarity (STS). Table 5 details corresponding instruction formats.
@@ 3.2.1 Bidirectional attention adaptation
For DMRetriever initialized from decoder-only backbones, we replace the causal attention mask with an all-ones matrix (B-1 in Figure 2) to enable the model to capture information from future tokens, inspired by the recent success of Lee et al. (2024a); BehnamGhader et al. (2024); Muennighoff et al. (2024). This replacement allows each token to attend to every other token in the sequence, converting it into a bidirectional model.
To make model aware of its bidirectional attention, we train model by using masked language model (MLM) training objective (Devlin et al., 2019). Given an input sequence, we randomly mask some of tokens and then train model to predict masked tokens based on past and future tokens.
@@ 3.2.2 MLM training data construction
We perform MLM training using a mix of general-domain and disaster management corpora. For the general domain, we use the English Wikipedia from Merity et al. (2016). For the disaster management domain, we use passages from DM-MTP (constructed in §3.3.1).
@@ 3.3.1 Pre-training dataset construction
In the absence of paired datasets for disaster management, DM-MTP represents the first large-scale collection of query–passage pairs in this domain. Inspired by Wang et al. (2023) and Yin et al. (2025), we propose a three-stage pipeline to generate it (A-1 in Figure 2, lines 1–7 in Algorithm 2). In the first stage, we crawl the web with queries formed by combining disaster event type names and location names, collect domain-specific PDFs, and process them into chunks, taking these as disaster-domain passages. Secondly, for each passage, a LLM generates information-need statement taking passage content and its designated search intent definition as input. Finally, the information-need statement is paired with the passage to prompt the LLM to generate a user query and a directly relevant passage, which serves as the positive passage for this query in DM-MTP. Appendix A details the whole DM-MTP construction process.
@@
For negative sampling in MTP, instead of explicitly mining hard negatives, we adopt the in-batch negative strategy (Karpukhin et al., 2020), where passages from other pairs within the same mini-batch serve as negative examples.
@@ 3.3.2 Contrastive Pre-training
DMRetriever is pre-trained on MTP using contrastive learning to separate relevant query–passage pairs from irrelevant ones. For each mini-batch $\mathcal{B}$, we adopt InfoNCE loss (Chen et al., 2020) as pre-training objective, encouraging similarity between $q_{i}$ and its positive passage $p_{i}^{+}$ to be higher than that with any other passage $p_{j}\in\mathcal{B},\,j\neq i$ within the same batch (B-2 in Figure 2):
@@ 3.4.2 Mutual-agreement False Positive Filter
Previous studies (Alberti et al., 2019; Xu et al., 2024b; Wang et al., 2022) show that LLM synthetic data can be noisy, and they commonly adopt a consistency-based filtering (CBF for short) approach: discarding pairs in which LLM-generated positive passage is not ranked within the top-$k$ results returned by a reference embedding model.
@@ 3.4.3 Difficulty-aware Hard Negative Mining
Inspired by Moreira et al. (2024), for a query $q_{i}$ with a positive passage $p_{i}^{+}$, a passage $p_{k}$ is treated as a hard negative $p_{i,k}^{-}$ if
@@
thenlper-gte-small (Li et al., 2023) 33M 18.04 9.13 10.95 49.63 37.51 55.55 30.14
arctic-embed-m (Merrick et al., 2024) 109M 33.15 14.04 8.48 35.07 38.67 56.20 30.94
thenlper-gte-base (Li et al., 2023) 109M 9.18 5.42 37.91 60.45 42.52 46.07 33.59
arctic-embed-m-v1.5 (Merrick et al., 2024) 109M 25.76 30.41 17.95 47.97 42.88 64.16 38.19
arctic-embed-s (Merrick et al., 2024) 33M 38.58 28.81 21.33 47.21 39.85 66.96 40.46
bge-small-en-v1.5 (Xiao et al., 2024) 33M 56.91 51.19 25.15 55.17 32.87 64.54 47.64
bge-base-en-v1.5 (Xiao et al., 2024) 109M 51.50 52.78 46.72 59.93 41.16 68.63 53.45
@@
arctic-embed-m-long (Merrick et al., 2024) 137M 21.51 10.86 19.24 36.13 41.67 54.94 30.73
arctic-embed-l (Merrick et al., 2024) 335M 40.56 30.19 14.98 32.64 34.20 56.10 34.78
bge-large-en-v1.5 (Xiao et al., 2024) 335M 56.76 54.45 32.20 54.90 35.11 64.47 49.65
gte-base-en-v1.5 (Li et al., 2023) 137M 60.51 55.62 46.26 52.24 39.59 70.40 54.10
mxbai-embed-large-v1 (Lee et al., 2024b) 335M 64.24 62.63 39.94 58.12 40.18 68.01 55.52
arctic-embed-m-v2.0 (Merrick et al., 2024) 305M 61.22 62.20 47.01 57.79 42.29 64.51 55.84
@@
arctic-embed-l-v2.0 (Merrick et al., 2024) 568M 55.23 59.11 38.11 60.10 41.07 62.61 52.70
gte-large-en-v1.5 (Li et al., 2023) 434M 67.37 58.18 39.43 52.66 34.45 66.47 53.09
Qwen3-Embedding-0.6B (Zhang et al., 2025) 596M 66.10 52.31 62.38 64.89 50.30 67.39 60.56
mulling-e5-large-instruct (Wang et al., 2024) 560M 67.97 64.64 62.25 66.78 48.51 63.42 62.26
mulling-e5-large (Wang et al., 2024) 560M 66.99 64.01 62.81 59.87 50.93 74.12 63.12
gte-Qwen2-1.5B-instruct (Li et al., 2023) 1.5B 69.85 59.17 65.09 62.73 55.51 73.58 64.32
inf-retriever-v1-1.5b (Junhan Yang, 2025) 1.5B 69.41 64.29 62.99 65.39 54.03 73.92 65.01
@@
Qwen3-Embedding-8B (Zhang et al., 2025) 7.6B 44.21 34.38 41.56 42.04 32.53 42.95 39.61
gte-Qwen2-7B-instruct (Li et al., 2023) 7.6B 70.24 47.41 63.08 31.62 53.71 74.88 56.82
NV-Embed-v1 (Moreira et al., 2024) 7.9B 68.06 62.70 56.02 59.64 48.05 67.06 60.26
Qwen3-Embedding-4B (Zhang et al., 2025) 4B 67.20 59.14 65.28 67.16 53.61 58.51 61.82
e5-mistral-7b-instruct (Wang et al., 2023) 7.1B 65.57 64.97 63.31 67.86 47.55 66.48 62.58
NV-Embed-v2 (Lee et al., 2024a) 7.9B 74.47 69.37 42.40 68.32 58.20 76.07 64.80
inf-retriever-v1 (Junhan Yang, 2025) 7.1B 72.84 66.74 66.23 65.53 51.86 75.98 66.53
SFR-Embedding-Mistral (Rui Meng, 2024) 7.1B 71.41 67.14 69.45 70.31 50.93 72.67 66.99
Linq-Embed-Mistral (Junseong Kim, 2024) 7.1B 74.40 70.31 64.11 70.64 52.46 71.25 67.19
@@ 3.4.4 Progressive Supervised Instruction Fine-tuning
Previous studies show that retrieval performance degrades when trained with negatives that are either too difficult or too simple (Merrick et al., 2024). To address this, for small-sized variants of DMRetriever (33M, 109M, and 335M) that have limited learning capacity after pre-training, we introduce progressive instruction fine-tuning (B-3 in Figure 2, Lines 9–13 in Algorithm 1), inspired by curriculum learning (Bengio et al., 2009). This approach gradually increases the difficulty of training samples across iterations, [^3] An iteration refers to a curriculum stage in Lines 9–12 of Algorithm 1, each consisting of multiple training batches. where the best validation checkpoint from one iteration initializes the next. At each stage, the difficulty level $\alpha$ is raised to make negatives harder, with models progressively trained on MTT-0.65, 0.75, and 0.85.

### v2

@@ 1 Introduction
Disasters, both natural and man-made, inevitably disrupt communities, endanger lives, and cause severe economic losses (fan2021disaster; liu2025artificial). During such events, emergency responders, policymakers, and affected populations rely on information retrieval (IR) systems for decision-making, situational awareness, and coordination of relief efforts (abbas2025exploring). Timely and reliable information access is thus essential for effective disaster management (priya2020taqe).
@@
Information needs during disaster events are highly diverse, spanning fact-checking critical information, monitoring social media, and consulting technical or policy documents (See §3.1 for full lists of search intents) (yin2025disastir). However, a recent analysis of 30 popular general-domain IR models finds that none consistently achieve SOTA performance across all disaster management-specific search intents (yin2025disastir). This limits their real-world applicability: some queries yield valuable insights, while others return irrelevant results, delaying decision-making and risking missed evidence that could compromise response efforts.
@@
Figure 2: Overview of data refinement pipeline and progressive fine-tuning for DMRetriever. MTP: Massive Text Pairs, MTT: Massive Text Triplets. Both MTP and MTT in sub-figure (B) include DM (Disaster Management) and GD (General Domain) variants, with constructions of GD variants omitted here and detailed in §3.3.1 and §3.4.3. DM passages include those extracted from PDFs and those generated by LLMs. DMRetriever-Init: backbone after bidirectional attention adaptation for decoder-only; original backbone for encoder-only. DMRetriever-PT: model after pre-training. DMRetriever-It2: model after iteration 2. Algorithm 1 details the full workflow.
@@
The success of domain-specific IR models in fields such as biomedicine (xu2024bmretriever), law (chalkidis2020legal), and finance (sarmah2024hybridrag) shows the effectiveness of adding deep domain knowledge to improve performance. However, injecting disaster management knowledge into DMRetriever to improve performance poses two challenges: (1) the lack of high-quality, knowledge-rich data that covers diverse search intents during disasters, and (2) the absence of an efficient training method that can transfer such knowledge across models of different sizes.
To tackle the first challenge, we propose a data refinement pipeline (A in Figure 2) to generate the first, knowledge-rich data for disaster management. We first prompt a large language model (LLM) [^1] We use GPT-4o-mini as LLM in this work. with domain-specific passages to generate large-scale unlabeled query–passage pairs [^2] Following xiao2024c; wang2022text, we treat data without mined negatives as unlabeled and data with mined hard negatives as labeled. (A-1 in Figure 2, §3.3.1). Given limitations of existing IR models in disaster management, we refine these noisy corpora with mutual-agreement filtering, which removes mislabeled pairs through model consensus (A-2 in Figure 2, §3.4.2). Recognizing the varied learning capacities of different-sized DMRetriever models, we further propose difficulty-aware hard negative mining to generate negatives at varying difficulty levels (A-3 in Figure 2, §3.4.3), finally yielding a high-quality labeled dataset.
For the second challenge, we propose a Progressive instruction fine-tuning (B in Figure 2) that enables DMRetriever models of different sizes to effectively learn domain knowledge. For large variants initialized from decoder-only backbones (Table 1), we adapt their causal attention into bidirectional attention, allowing access to information from future tokens (B-1 in Figure 2, §3.2). Next, to better initialize DMRetriever and inject domain knowledge, we perform large-scale contrastive pre-training on unlabeled query–passage pairs (B-2 in Figure 2, §3.3). Finally, to adapt DMRetriever to diverse search intents, we perform instruction fine-tuning on labeled data2, where for small variants that have limited learning capacity, the difficulty of negative samples is progressively increased during fine-tuning (B-3 in Figure 2, §3.4.4).
@@
    We introduce progressive instruction fine-tuning which enables models of varied sizes to effectively absorb domain knowledge and enhances their ability across different search intents.
@@ 2 Related Work
General-domain retrieval has been long dominated by bidirectional encoders such as BERT (devlin2019bert) and T5 (raffel2020exploring), with models like Sentence-BERT (reimers2019sentence) and SimCSE (gao2021simcse) improving performance through contrastive pre-training and supervised fine-tuning on MS MARCO (nguyen2016ms). More recently, decoder-only models, from GPT-based embeddings (brown2020language) to models such as Qwen3-Embedding (qwen3embedding), have leveraged instruction tuning and large-scale synthetic data to achieve SOTA results.
However, these general-domain models face clear limitations in disaster management, with no approach achieving SOTA performance across diverse search intents (yin2025disastir) due to distribution shift issue (thakur2021beir). Prior work on disaster information access has mainly targeted on social media, emphasizing clustering of posts, event or situational classification, and named entity recognition (imran2015processing; alam2021crisisbench; yin2024crisissense). While valuable, these efforts remain narrow in scope and have yet to address the need for robust retrieval systems tailored to disaster management-related diver search intents.
@@ 3.1 Preliminaries
In DMRetriever, queries and passages are encoded into dense vectors using a shared embedding model $\mathbf{E}(\cdot)$. Given a query $q$ and a passage $p$, we prepend a search intent-specific instruction $I_{q}$ to the query to improve search intent awareness (wei2021finetuned; ouyang2022training; lee2024nv).
Following DisastIR (yin2025disastir), we consider six disaster management-specific search intents: question-answer (QA), Twitter (TW), Fact Checking (FC), Natural Language Inference (NLI), and Semantic Textual Similarity (STS). Table 5 details corresponding instruction formats.
@@ 3.2.1 Bidirectional attention adaptation
For DMRetriever initialized from decoder-only backbones, we replace the causal attention mask with an all-ones matrix (B-1 in Figure 2) to enable the model to capture information from future tokens, inspired by the recent success of lee2024nv; behnamghader2024llm2vec; muennighoff2024generative. This replacement allows each token to attend to every other token in the sequence, converting it into a bidirectional model.
To make model aware of its bidirectional attention, we train model by using masked language model (MLM) training objective (devlin2019bert). Given an input sequence, we randomly mask some of tokens and then train model to predict masked tokens based on past and future tokens.
@@ 3.2.2 MLM training data construction
We perform MLM training using a mix of general-domain and disaster management corpora. For the general domain, we use the English Wikipedia from merity2016pointer. For the disaster management domain, we use passages from DM-MTP (constructed in §3.3.1).
@@ 3.3.1 Pre-training dataset construction
In the absence of paired datasets for disaster management, DM-MTP represents the first large-scale collection of query–passage pairs in this domain. Inspired by wang2023improving and yin2025disastir, we propose a three-stage pipeline to generate it (A-1 in Figure 2, lines 1–7 in Algorithm 2). In the first stage, we crawl the web with queries formed by combining disaster event type names and location names, collect domain-specific PDFs, and process them into chunks, taking these as disaster-domain passages. Secondly, for each passage, a LLM generates information-need statement taking passage content and its designated search intent definition as input. Finally, the information-need statement is paired with the passage to prompt the LLM to generate a user query and a directly relevant passage, which serves as the positive passage for this query in DM-MTP. Appendix A details the whole DM-MTP construction process.
@@
For negative sampling in MTP, instead of explicitly mining hard negatives, we adopt the in-batch negative strategy (karpukhin2020dense), where passages from other pairs within the same mini-batch serve as negative examples.
@@ 3.3.2 Contrastive Pre-training
DMRetriever is pre-trained on MTP using contrastive learning to separate relevant query–passage pairs from irrelevant ones. For each mini-batch $\mathcal{B}$, we adopt InfoNCE loss (chen2020simple) as pre-training objective, encouraging similarity between $q_{i}$ and its positive passage $p_{i}^{+}$ to be higher than that with any other passage $p_{j}\in\mathcal{B},\,j\neq i$ within the same batch (B-2 in Figure 2):
@@ 3.4.2 Mutual-agreement False Positive Filter
Previous studies (alberti2019synthetic; xu2024bmretriever; wang2022text) show that LLM synthetic data can be noisy, and they commonly adopt a consistency-based filtering (CBF for short) approach: discarding pairs in which LLM-generated positive passage is not ranked within the top-$k$ results returned by a reference embedding model.
@@ 3.4.3 Difficulty-aware Hard Negative Mining
Inspired by moreira2024nv, for a query $q_{i}$ with a positive passage $p_{i}^{+}$, a passage $p_{k}$ is treated as a hard negative $p_{i,k}^{-}$ if
@@
thenlper-gte-small (li2023gte) 33M 18.04 9.13 10.95 49.63 37.51 55.55 30.14
arctic-embed-m (merrick2024arctic) 109M 33.15 14.04 8.48 35.07 38.67 56.20 30.94
thenlper-gte-base (li2023gte) 109M 9.18 5.42 37.91 60.45 42.52 46.07 33.59
arctic-embed-m-v1.5 (merrick2024arctic) 109M 25.76 30.41 17.95 47.97 42.88 64.16 38.19
arctic-embed-s (merrick2024arctic) 33M 38.58 28.81 21.33 47.21 39.85 66.96 40.46
bge-small-en-v1.5 (xiao2024c) 33M 56.91 51.19 25.15 55.17 32.87 64.54 47.64
bge-base-en-v1.5 (xiao2024c) 109M 51.50 52.78 46.72 59.93 41.16 68.63 53.45
@@
arctic-embed-m-long (merrick2024arctic) 137M 21.51 10.86 19.24 36.13 41.67 54.94 30.73
arctic-embed-l (merrick2024arctic) 335M 40.56 30.19 14.98 32.64 34.20 56.10 34.78
bge-large-en-v1.5 (xiao2024c) 335M 56.76 54.45 32.20 54.90 35.11 64.47 49.65
gte-base-en-v1.5 (li2023gte) 137M 60.51 55.62 46.26 52.24 39.59 70.40 54.10
mxbai-embed-large-v1 (emb2024mxbai) 335M 64.24 62.63 39.94 58.12 40.18 68.01 55.52
arctic-embed-m-v2.0 (merrick2024arctic) 305M 61.22 62.20 47.01 57.79 42.29 64.51 55.84
@@
arctic-embed-l-v2.0 (merrick2024arctic) 568M 55.23 59.11 38.11 60.10 41.07 62.61 52.70
gte-large-en-v1.5 (li2023gte) 434M 67.37 58.18 39.43 52.66 34.45 66.47 53.09
Qwen3-Embedding-0.6B (qwen3embedding) 596M 66.10 52.31 62.38 64.89 50.30 67.39 60.56
mulling-e5-large-instruct (wang2024multilingual) 560M 67.97 64.64 62.25 66.78 48.51 63.42 62.26
mulling-e5-large (wang2024multilingual) 560M 66.99 64.01 62.81 59.87 50.93 74.12 63.12
gte-Qwen2-1.5B-instruct (li2023gte) 1.5B 69.85 59.17 65.09 62.73 55.51 73.58 64.32
inf-retriever-v1-1.5b (infly-ai_2025) 1.5B 69.41 64.29 62.99 65.39 54.03 73.92 65.01
@@
Qwen3-Embedding-8B (qwen3embedding) 7.6B 44.21 34.38 41.56 42.04 32.53 42.95 39.61
gte-Qwen2-7B-instruct (li2023gte) 7.6B 70.24 47.41 63.08 31.62 53.71 74.88 56.82
NV-Embed-v1 (moreira2024nv) 7.9B 68.06 62.70 56.02 59.64 48.05 67.06 60.26
Qwen3-Embedding-4B (qwen3embedding) 4B 67.20 59.14 65.28 67.16 53.61 58.51 61.82
e5-mistral-7b-instruct (wang2023improving) 7.1B 65.57 64.97 63.31 67.86 47.55 66.48 62.58
NV-Embed-v2 (lee2024nv) 7.9B 74.47 69.37 42.40 68.32 58.20 76.07 64.80
inf-retriever-v1 (infly-ai_2025) 7.1B 72.84 66.74 66.23 65.53 51.86 75.98 66.53
SFR-Embedding-Mistral (SFRAIResearch2024) 7.1B 71.41 67.14 69.45 70.31 50.93 72.67 66.99
Linq-Embed-Mistral (LinqAIResearch2024) 7.1B 74.40 70.31 64.11 70.64 52.46 71.25 67.19
@@ 3.4.4 Progressive Supervised Instruction Fine-tuning
Previous studies show that retrieval performance degrades when trained with negatives that are either too difficult or too simple (merrick2024arctic). To address this, for small-sized variants of DMRetriever (33M, 109M, and 335M) that have limited learning capacity after pre-training, we introduce progressive instruction fine-tuning (B-3 in Figure 2, Lines 9–13 in Algorithm 1), inspired by curriculum learning (bengio2009curriculum). This approach gradually increases the difficulty of training samples across iterations, [^3] An iteration refers to a curriculum stage in Lines 9–12 of Algorithm 1, each consisting of multiple training batches. where the best validation checkpoint from one iteration initializes the next. At each stage, the difficulty level $\alpha$ is raised to make negatives harder, with models progressively trained on MTT-0.65, 0.75, and 0.85.